1.determine whether the given line and the given planes are parallel:

1.1.x=1+t,y=-1-t,z=-2t and x+2y+3z-9=0.

1.2.<0,1,2>+t<3,2,-1> and 4x-y+2z+1=0.


2.

2.1.find parametric equations of the line that passes through the point p=(2,0,-1) and is parallel to the vector n=<2,1,3>.

2.2.find parametric equations of the line that passes through the points a=(1,2,-3) and b=(7,2,-4).

2.3.find parametric equations for the line intersection of the planes -5x+y-2z=3 and 2x-3y+5z=-7.


1
Expert's answer
2021-06-30T16:10:01-0400

1.1. The plane "x+2y+3z-9=0" is orthogonal to the vector "\vec{n}=(1,2,3)".

The line "x=1+t,\,y=-1-t,\,z=-2t" has a direction vector "\vec{v}=(1,-1,-2)".

It is parallel to the plane with a normal vector "\vec{n}=(1,2,3)" , if and only if "\langle \vec{v}, \vec{n}\rangle=0".

But since "\langle \vec{v}, \vec{n}\rangle = 1\cdot 1+(-1)\cdot 2+(-2)\cdot 3=-8\ne 0", this line and this plane are not parallel.


1.2. The plane "4x-y+2z+1=0" is orthogonal to the vector "\vec{n}=(4,-1,2)".

The line "<0,1,2>+t<3,2,-1>" has a direction vector "\vec{v}=(3,2,-1)".

It is parallel to the plane with a normal vector "\vec{n}=(4,-1,2)" , if and only if "\langle \vec{v}, \vec{n}\rangle=0".

But since "\langle \vec{v}, \vec{n}\rangle = 3\cdot 4+2\cdot (-1)+(-1)\cdot 2=8\ne 0", this line and this plane are not parallel.


2.1.find parametric equations of the line that passes through the point p=(2,0,-1) and is parallel to the vector n=<2,1,3>.

"x(t)=2+2t,\, y(t)=t,\, z(t) = -1+3t"


2.2.find parametric equations of the line that passes through the points a=(1,2,-3) and b=(7,2,-4).

"\vec{ab}=(7,2,-4)-(1,2,-3)=(6,0,-1)"

"x(t)=1+6t,\, y(t)=2,\, z(t) = -3-t"


2.3.find parametric equations for the line intersection of the planes -5x+y-2z=3 and 2x-3y+5z=-7.

Let z=t be a parameter. Then

-5x+y=3+2z=3+2t, y=5x+3+2t

2x-3y=-5z-7, 2x-(5x+3+2t)=-5t-7

-3x=-3t-4

x=t+4/3, y=5(t+4/3)+3+2t = 7t+29/3, z=t. This is parametric equation for the line intersection of the given planes.